OpenStudy (anonymous):

A line passes through the point (4, 2) and has a slope of -5/4. Write an equation in slope-intercept form for this line.

OpenStudy (anonymous):

y=mx+b 2=(-5/4)(-4)+b 2=5+b b=-3 So the equation is y=(-5/4)x-3

OpenStudy (anonymous):

thank you!